# Microlensing optical depth, event rate, and limits on compact objects in dark matter based on 20 years of OGLE observations of the Small Magellanic Cloud

### v1

@@ 10.3 Summary
This paper concludes the current searches for microlensing events toward the Magellanic Clouds using data from the OGLE survey. In Paper 1 and Paper 2, we presented the results of the searches for very long-timescale microlensing events toward the LMC using the combined OGLE-III and OGLE-IV data. Such long-duration events were expected if dark matter in the Milky Way halo consisted of massive PBHs, with masses ranging from dozens to hundreds $M_{\odot}$, similar to black holes detected by the gravitational wave observatories. In Paper 3, we searched for short-duration microlensing events that could have been caused by planetary-mass PBHs. No such events were discovered, which provided very strict limits on the abundance of PBHs and other compact objects in the Milky Way halo with masses ranging from $10^{-8}$ to $10^{3}\,M_{\odot}$.
@@
We recovered three previously known events and discovered three new ones. We measured the microlensing optical depth toward the SMC to be $\tau=(0.32\pm 0.18)\times 10^{-7}$ and the event rate to be $\Gamma=(1.18\pm 0.57)\times 10^{-7}\,\mathrm{yr}^{-1}\,\mathrm{star}^{-1}$. The properties of the detected events are consistent with lenses originating from known stellar populations within the SMC or in the Milky Way disk, leaving little room to PBHs or other dark compact objects in the dark matter halo.
@@
The data presented in this paper are publicly available at https://www.astrouw.edu.pl/ogle/ogle4/SMC_OPTICAL_DEPTH/.
@@ Acknowledgements
We thank all the OGLE observers for their contribution to the collection of the photometric data over the decades. We thank Łukasz Wyrzykowski for his comments on the manuscript. This research was funded in part by National Science Centre, Poland, grants OPUS 2021/41/B/ST9/00252 and SONATA 2023/51/D/ST9/00187 awarded to P.M. The OGLE project has received funding from the Polish National Science Centre grant OPUS 2024/55/B/ST9/00447 awarded to A.U.
@@ Appendix A SMC Proper Motion and Rotation
Table 9: Parameters of the SMC Proper Motion Model

### v2

@@ 10.3 Summary
This paper concludes the current searches for microlensing events toward the Magellanic Clouds using data from the OGLE survey. In Paper 1 and Paper 2, we presented the results of the searches for very long-timescale microlensing events toward the LMC using the combined OGLE-III and OGLE-IV data. Such long-duration events were expected if dark matter in the Milky Way halo consisted of massive PBHs, with masses ranging from dozens to hundreds $M_{\odot}$, similar to black holes detected by the gravitational-wave observatories. In Paper 3, we searched for short-duration microlensing events that could have been caused by planetary-mass PBHs. No such events were discovered, which provided very strict limits on the abundance of PBHs and other compact objects in the Milky Way halo with masses ranging from $10^{-8}$ to $10^{3}\,M_{\odot}$.
@@
We recovered three previously known events and discovered three new ones. We measured the microlensing optical depth toward the SMC to be $\tau=(0.32\pm 0.18)\times 10^{-7}$ and the event rate to be $\Gamma=(1.18\pm 0.57)\times 10^{-7}\,\mathrm{yr}^{-1}\,\mathrm{star}^{-1}$. The properties of the detected events are consistent with lenses originating from known stellar populations within the SMC or in the Milky Way disk, leaving little room for PBHs or other dark compact objects in the dark matter halo.
@@
The data presented in this paper are publicly available at https://www.astrouw.edu.pl/ogle/ogle4/SMC_OPTICAL_DEPTH/ and via doi: 105281/zenodo.16735482 (catalog 105281/zenodo.16735482).
@@ Acknowledgements
We thank the anonymous referee for providing constructive comments that helped improve the paper. We thank all of the OGLE observers for their contribution to the collection of the photometric data over the decades. We thank Łukasz Wyrzykowski for his comments on the manuscript. This research was funded in part by National Science Centre, Poland, grants OPUS 2021/41/B/ST9/00252 and SONATA 2023/51/D/ST9/00187 awarded to P.M. The OGLE project has received funding from the Polish National Science Centre grant OPUS 2024/55/B/ST9/00447 awarded to A.U.
@@ Appendix A SMC Proper Motion and Rotation
Table 10: Parameters of the SMC Proper-Motion Model